The equation of the line that goes through the point (7,7) and is perpendicular to the line 2x+5y=4
Answer by Ave(106)

First put that equation in slope intercept so we can find the slope
5y=-2x+4
y=-2/5x+4/5
Slope is -2/5, so the slope of a perpendicular line would be 5/2
Now use point slope form to find the new equation
y-y1=slope(x-x1) x1, and y1 are the points from (7,7)
y-7=5/2(x-7) multiply both sides by 2 to get rid of fraction
2y-14=5x-35
2y=5x-21
----
y=5/2x-21/2<- Answer in slope intercept form
----
-5x+2y=-21 <- Answer in standard form